OpenStudy (anonymous):

what about this one 4=2+p/2

OpenStudy (anonymous):

p/2 = 2 p = 4

OpenStudy (anonymous):

first shift 2 to LHS 4-2 = p/2 2=p/2 then multiply both sides by 2 2*2 = p/2*2 4 = p